### v1

@@ Abstract
Context. This research presents unsupervised machine learning and statistical methods to identify and analyze tidal tails in open star clusters using data from the Gaia-DR3 catalog.
Aims. We aim to identify member stars, detect and analyze tidal tails in five open clusters, BH 164, Alessi 2, NGC 2281, NGC 2354, and M67, of ages between 60 Myr and 4 Gyr. These clusters were selected based on the previous evidence of extended tidal structures.
Methods. We utilized machine learning algorithms such as Density-Based Spatial Clustering of Applications with Noise (DBSCAN) and Principal Component Analysis (PCA), along with statistical methods to analyze the kinematic, photometric, and astrometric properties of stars. Key characteristics of tidal tails, including radial velocity, color-magnitude diagram, and spatial projections in the tangent plane beyond the cluster’s Jacobi radius ($r_{J}$), were used to detect them. We used N-body simulations to visualize and compare the observables with real data. Further analysis was done on the detected cluster and tail stars to study their internal dynamics and populations, including binary fraction. We also applied the residual velocity method to detect rotational patterns in the clusters and their tails.
Results. We identified tidal tails in all five clusters, with detected tails extending farther in some clusters and containing significantly more stars than previously reported (tails ranging from 40 to 100 pc, one to four times the Jacobi radius, with 100 - 200 tail stars). The luminosity functions of the tails and their parent clusters were generally consistent, and tails lacked massive stars. In general, the binary fraction was found to be higher in the tidal tails. Significant rotation was detected in M67 and NGC 2281 for the first time.
@@ 1 Introduction
The Milky Way, our home galaxy, is a barred spiral galaxy with a central bulge, spiral arms, and a thin and thick disc (Hou & Han, 2014; Shen & Zheng, 2020). This complicated structure also contains a dark matter halo, which affects the motions of stars and stellar clusters (Posti & Helmi, 2019; Deason et al., 2019). The Milky Way’s stellar population is diversified, with old, metal-poor stars in the halo and new, metal-rich stars in the thin disc. This complex stellar environment contains a variety of stellar groups, including open clusters, globular clusters, and tidal streams, each of which serves as a natural laboratory for studying galactic dynamics, star formation, and stellar evolution (Kharchenko et al., 2012, 2013; Vasiliev, 2019; Helmi, 2020; Cantat-Gaudin, 2022).
@@
The high stellar density, strong gravitational binding, and reduced field star contamination make the tidal tails of globular clusters easier to detect (Leon et al., 2000; Odenkirchen et al., 2001, 2003). The low stellar densities of open clusters and substantial field star contamination from the dense stellar background of the Galactic disc pose significant challenges for detecting the tidal tails of open clusters (Bergond et al., 2001; Chumak et al., 2010; Röser et al., 2011). Also, because of internal relaxation, interactions with molecular clouds, and the disruptive effects of the Galactic tidal field, open clusters frequently dissolve over time, leaving their tidal tails poorly populated and difficult to identify.
@@
In particular, Meingast et al. (2019) discovered the Pisces-Eridanus stream using wavelet decomposition and Density-Based Spatial Clustering of Applications with Noise (DBSCAN), identifying 256 stars in a 400 pc structure. Subsequent work by Ratzenböck et al. (2020) and Röser & Schilbach (2020) expanded membership using support vector machines and convergent point methods. Kounkel & Covey (2019) and Kounkel et al. (2020) applied hierarchical density-based clustering (HDBSCAN) to the Gaia DR2 and identified 1312 stellar groups and 328 stellar strings. Meingast et al. (2021) used Gaia DR2 to identify the tidal tails and coronae in ten open clusters and developed a velocity-based member identification method, starting with the Cantat-Gaudin et al. (2018) membership tables. (Gao, 2020a, b) applied Principal Component Analysis (PCA) and GMM to identify extra-tidal stars in M67 and NGC 2506, respectively, using Gaia DR2.
@@
Bhattacharya et al. (2022) employed ML-MOC, a membership algorithm based on k-nearest neighbors (kNN) and GMM for precise identification from Gaia EDR3 data, and identified 46 open clusters having a stellar corona, while Tarricq et al. (2022) used HDBSCAN to find members up to 50 pc from cluster centers, fitted King profiles, used GMMs to analyze spatial distributions, and reported the detection of 71 OCs with tidal tails. Zucker et al. (2022) used a combination of Gaia DR2 and Gaia DR3 to investigate the stellar strings identified in Kounkel & Covey (2019) by manually linking stellar groups, while Andrews et al. (2022) analyzed Theia 456 using Gaia EDR3 and produced a more robust membership catalog of 362 stars using DBSCAN. Jerabkova et al. (2021) used the compact convergent point method to recover 800 pc long tidal tails of the Hyades star cluster using Gaia DR2 and EDR3.
@@
Using Gaia DR3, Hunt & Reffert (2023) applied the HDBSCAN algorithm to recover 7167 clusters, also including tidal tails for some clusters. However, Hunt & Reffert (2024) derived completeness-corrected photometric masses for 6956 clusters from the earlier work and used these masses to compute their Jacobi radius. Kos (2024) simulated cluster dissolution to estimate the likelihood of star positions, comparing these with Gaia DR3 data to refine membership probabilities. Recently, Risbud et al. (2025) used the convergent point method to identify tidal tails in 19 open clusters within 400 pc.
@@
The previous methods often struggle with limitations such as poor astrometric precision, model dependencies, or applicability limited to nearby clusters. These constraints significantly reduce the number of clusters that can be studied in detail. To address this, we designed a method that uses optimized clustering and kinematic analysis to identify tidal tails even in distant systems. In this work, we have used the Gaia-DR3 dataset to identify and investigate tidal tails in five open clusters: BH 164, NGC 2354, Alessi 2, NGC 2281, and M67. We tuned the algorithms to increase the detected extent of tidal tails, incorporating radial velocity (RV) filtering (where available), color-magnitude diagram (CMD) matching for age and composition consistency, and an unsupervised approach using PCA to determine the principal axis of elongation, followed by DBSCAN for tail detection. The identified tails were validated using the Jacobi radius to confirm their tidal origin. Projection effects were taken into account in the tangential plane, providing a clearer spatial distribution of the tails. We detected a greater number of stars in the four clusters and identified longer tidal tails in three of them compared to previous studies. Additionally, we obtained a more accurate estimate of the projected tail extent for M67, whose orientation aligns well with the expected orbit, consistent with N-body simulations. Furthermore, rotation was studied in the M67 and NGC 2281 clusters and tails. The residual velocity method, adopted from Hao et al. (2022, 2024), was previously used exclusively for member stars to confirm internal rotation in clusters such as Praesepe, Pleiades, Alpha Persei, and Hyades. In this study, we applied it to the cluster and its tails to analyze their rotational patterns.
@@
Section 2 outlines the application of DBSCAN to identify the core cluster and also explains the N-body simulations used for reference. Section 3 presents the methods used to detect tidal tails. In Section 4, we present a further detailed analysis utilizing proper motion vector analysis, orbit integration, luminosity functions, and binary fractions to explore the cluster and tail dynamics. Section 5 presents our tidal tail findings and discusses the results. In Section 6, the rotational aspect of M67 and NGC 2281 is explored using the residual motion of stars to study the presence of three-dimensional rotation (Hao et al., 2022, 2024). Finally, we summarize and conclude our research in Section 7.
@@ 2.2 Gaia DR3
The latest Gaia Data Release 3 (DR3; Gaia Collaboration et al. 2023) provides a substantial update over Gaia DR2, contains a total of 34 months of observational data, and features major advancements in data processing techniques. Gaia DR3 catalogs 1.47 billion sources with 5- or 6-parameter astrometric solutions, achieving a 30% improvement in parallax precision and nearly doubling the accuracy of proper motion measurements. These enhancements significantly improve the visibility of open clusters (OCs) within the Gaia dataset, particularly in proper motion diagrams. The signal-to-noise ratio (S/N) for identifying distant OCs has increased by approximately four (Hunt & Reffert, 2023), benefiting from a more compact Gaussian distribution of cluster stars for those with proper motion dispersions now within the reduced Gaia error margins.
@@
  • BP–RP (in mag): colour index from Gaia blue and red photometers. Typical errors: $\lesssim$0.01 mag (bright), $\approx$0.05 mag (faint).
@@
  • radial_velocity (in km s-1): line-of-sight velocity, available for bright stars ($G\lesssim 15$). Typical errors: $\approx$0.2–5 km s-1.
Figure 1: The spatial distribution, proper motion, and parallax of the Filtered Sample (light coral) after filtering and the Cluster members (blue) after applying DBSCAN for M67.
@@ 2.3 Data Filtering and Preparation
  • Proper motion (PM) filtering: pmra_min $\leq$ pmra $\leq$ pmra_max and  pmdec_min $\leq$ pmdec $\leq$ pmdec_max.
    For example, for M67, the ranges used were: $12.5\,\text{mas yr${}^{-1}$}\leq\texttt{pmra}\leq-9.5\,\text{mas yr${}^{-1}$},\quad-4\,\text{mas yr${}^{-1}$}\leq\texttt{pmdec}\leq-1\,\text{mas yr${}^{-1}$}$. These bounds were chosen based on the literature values (Gao, 2020a) and by visually inspecting the star distribution in proper motion space.
  • Error Filtering: Stars with proper motion errors <0.5<0.5 mas yr-1 were retained. We limited the parallax ($\varpi$) using $\varpi>0$ and $\sigma_{\varpi}/\varpi<0.1$. Stars passing these filters were labeled as the Filtered Sample. We used relative parallax error to maintain consistent distance precision across clusters, although this imposes stricter absolute error thresholds on distant clusters, potentially biasing against the inclusion of faint members. Additionally, unresolved binaries may introduce astrometric errors not explicitly accounted for in this analysis. The spatial, PM, and parallax distributions for M67 are highlighted in light coral in Figure 1.
  • Data Preparation: For core cluster identification, we used the astrometric features: ra, dec, pmra, pmdec, and distance. We used distance instead of parallax to linearize spatial separations, enabling meaningful Euclidean distances in the clustering feature space. Since we filtered stars to have relative parallax error <0.1<0.1, these distances are considered reliable. Stars with missing (NaN) values in any of these features were excluded. These core features were then normalized to zero mean and unit variance to ensure each contributed equally to the clustering process.
@@ 2.4 DBSCAN Clustering for Core Cluster Identification
Density-Based Spatial Clustering of Applications with Noise (DBSCAN) (Ester et al., 1996; Pedregosa et al., 2011) is a density-based clustering algorithm that identifies groups of closely packed points while treating outliers as noise. Unlike methods like $k$-means, DBSCAN does not require specifying the number of clusters in advance and can discover clusters of arbitrary shape, making it well-suited for astrophysical data where the boundaries of a star cluster may not be sharply defined. It gives clear-cut clusters without soft membership boundaries.
For a data point to be considered as a part of a cluster, it must have a minimum number of neighboring data points within a specified distance, $\epsilon$.
Key Parameters of DBSCAN:
@@
The DBSCAN algorithm identifies core points, border points, and noise points. Core points have sufficient neighbors within $\epsilon$; border points are neighbors of core points but are not core themselves; noise points are neither. In this work, DBSCAN was applied to the scaled 5D feature (ra, dec, pmra, pmdec, distance) space of the filtered sample to identify a dense overdensity corresponding to the core cluster population. For each cluster, the parameters ($\epsilon$, min_neighbours) were tuned to find the combination that resulted in the most spatially complete and centrally concentrated core in ra–dec space, while minimizing contamination.
In M67, this method isolated 1037 Cluster members (Nmemb), whose spatial, PM, and $\varpi$ distributions are also shown in blue in Figure 1.
@@ 3 Tidal Tail Detection
Once core members were identified, we searched for tidal tail stars among non-cluster Filtered Sample stars within 500 pc of the cluster’s mean distance. The approach combined radial velocity filtering, CMD matching, and machine learning with scikit-learn (Pedregosa et al., 2011), including KD-Tree (k-Dimensional Tree), PCA, and DBSCAN.
Tidal stars are expected to share radial velocities and CMD features with the cluster, align along the cluster’s elongation axis, and lie beyond the Jacobi radius (Meingast et al., 2021).
Table 1: Core cluster parameters of selected open clusters, including Column 1: Name; Columns 2–3: Equatorial coordinates; Columns 4–5: Galactic Lon-Lat; Column 6: Age from Bossini et al. (2019) and Cantat-Gaudin et al. (2020); Columns 7-10: Proper motions, Mean Parallax and Mean Distance of identified members; Column 11: Mean RV of identified core cluster members; Column 12: Number of identified core cluster members.
Name ra dec l b Age pmra pmdec $\varpi$ Dist RV $N_{memb}$
@@ 3.1 Radial Velocity Filtering
After defining the core cluster, we applied a radial velocity filter to further refine the potential tail candidates. Specifically, we selected stars within the 500 pc distance range whose Gaia radial velocities fall within a range of $\pm 10$ km s-1 of the core cluster’s mean radial velocity:
$\texttt{radial\_velocity}_{\text{tail}}=\texttt{radial\_velocity}_{\text{core}}\pm 10\,\text{km s${}^{-1}$}$ (1)
This $\pm 10$ km s-1 threshold serves as a loose filtering criterion to remove clear RV outliers that are inconsistent with the cluster’s dynamics. It is not based on an assumption of uniform radial velocity along the tail. Rather, it is broad enough to account for small projection-induced velocity gradients  [^1] Projection effects are corrected after the tidal tail detection for further analysis; see Section 4.1., which arise from geometric effects over large angular separations. This window was chosen to preserve plausible tail candidates while excluding clear outliers. Since Gaia DR3 provides RVs only for a subset of stars, this filtering step is applied only when radial_velocity data are available. For stars lacking RV data, we rely on CMD-based filtering to evaluate their consistency with the cluster’s stellar population.
3.2 Color-Magnitude Diagram (CMD) Matching with KDTree
In addition to radial velocity, we used Color (BP-RP) versus the apparent G Magnitude CMD to refine our identification of tidal tail stars. In the CMD space, tidal tail stars are expected to closely resemble the core cluster members, sharing similar age, metallicity, and evolutionary history. The CMD provides a visual representation of these properties and allows us to distinguish between stars that formed together and those that are unrelated to background or foreground stars.
To efficiently process large datasets like Gaia DR3, we employed a KD-Tree structure for performing the CMD cut. This algorithm enables fast nearest-neighbor searches in high-dimensional space and was used to identify stars whose CMD values closely match those of the core cluster members. The tree partitions the data recursively by splitting on dimensions like color and magnitude. A distance threshold of 0.05 in CMD space was applied to select potential tidal tail members, ensuring photometric similarity with the cluster core. The CMD distance is defined as:
$d_{\text{CMD}}=\sqrt{(G-G_{\text{core}})^{2}+\left[(BP-RP)-(BP_{\text{core}}-RP_{\text{core}})\right]^{2}}$ (2)

### v3

@@ Abstract
Context. This research presents unsupervised machine learning and statistical methods to identify and analyze tidal tails in open star clusters using data from the Gaia DR3 catalog.
Aims. We aim to identify member stars and to detect and analyze tidal tails in five open clusters, BH 164, Alessi 2, NGC 2281, NGC 2354, and M67, of ages between 60 Myr and 4 Gyr. These clusters were selected based on the previous evidence of extended tidal structures.
Methods. We utilized machine learning algorithms such as Density-based Spatial Clustering of Applications with Noise (DBSCAN) and principal component analysis (PCA), along with statistical methods to analyze the kinematic, photometric, and astrometric properties of stars. Key characteristics of tidal tails, including radial velocity, the color-magnitude diagram, and spatial projections in the tangent plane beyond the cluster’s Jacobi radius ($r_{J}$), were used to detect them. We used N-body simulations to visualize and compare the observables with real data. Further analysis was done on the detected cluster and tail stars to study their internal dynamics and populations, including the binary fraction. We also applied the residual velocity method to detect rotational patterns in the clusters and their tails.
Results. We identified tidal tails in all five clusters, with detected tails extending farther in some clusters and containing significantly more stars than previously reported (tails ranging from 40 to 100 pc, one to four times their $r_{J}$, with 100-200 tail stars). The luminosity functions of the tails and their parent clusters were generally consistent, and tails lacked massive stars. In general, the binary fraction was found to be higher in the tidal tails. Significant rotation was detected in M67 and NGC 2281 for the first time.
@@ 1 Introduction
The Milky Way, our home galaxy, is a barred spiral galaxy with a central bulge, spiral arms, and a thin and thick disk (Hou & Han, 2014; Shen & Zheng, 2020). This complicated structure also contains a dark matter halo, which affects the motions of stars and stellar clusters (Posti & Helmi, 2019; Deason et al., 2019). The Milky Way’s stellar population is diversified, with old, metal-poor stars in the halo and new, metal-rich stars in the thin disk. This complex stellar environment contains a variety of stellar groups, including open clusters, globular clusters, and tidal streams, with each of which serving as a natural laboratory for studying Galactic dynamics, star formation, and stellar evolution (Kharchenko et al., 2012, 2013; Vasiliev, 2019; Helmi, 2020; Cantat-Gaudin, 2022).
@@
The high stellar density, strong gravitational binding, and reduced field star contamination make the tidal tails of globular clusters easier to detect (Leon et al., 2000; Odenkirchen et al., 2001, 2003). The low stellar densities of open clusters and substantial field star contamination from the dense stellar background of the Galactic disk pose significant challenges for detecting the tidal tails of open clusters (Bergond et al., 2001; Chumak et al., 2010; Röser et al., 2011). Also, because of internal relaxation, interactions with molecular clouds, and the disruptive effects of the Galactic tidal field, open clusters frequently dissolve over time, leaving their tidal tails poorly populated and difficult to identify.
@@
In particular, Meingast et al. (2019) discovered the Pisces-Eridanus stream using wavelet decomposition and Density-Based Spatial Clustering of Applications with Noise (DBSCAN), identifying 256 stars in a 400 pc structure. Subsequent work by Ratzenböck et al. (2020) and Röser & Schilbach (2020) expanded membership using support vector machines and convergent point methods. Kounkel & Covey (2019) and Kounkel et al. (2020) applied hierarchical density-based clustering (HDBSCAN) to Gaia DR2 and identified 1312 stellar groups and 328 stellar strings. Meingast et al. (2021) used Gaia DR2 to identify the tidal tails and coronae in ten open clusters and developed a velocity-based member identification method, starting with the Cantat-Gaudin et al. (2018) membership tables. Gao (2020a, b) applied principal component analysis (PCA) and Gaussian mixture model (GMM) to identify extra-tidal stars in M67 and NGC 2506, respectively, using Gaia DR2.
@@
Bhattacharya et al. (2022) employed ML-MOC, a membership algorithm based on k-nearest neighbors (kNN) and GMM for precise identification from Gaia EDR3 data, and identified 46 open clusters with a stellar corona, while Tarricq et al. (2022) used HDBSCAN to find members up to 50 pc from cluster centers, fitted King profiles, used GMMs to analyze spatial distributions, and reported the detection of 71 open clusters with tidal tails. Zucker et al. (2022) used a combination of Gaia DR2 and Gaia DR3 to investigate the stellar strings identified in Kounkel & Covey (2019) by manually linking stellar groups, while Andrews et al. (2022) analyzed Theia 456 using Gaia EDR3 and produced a more robust membership catalog of 362 stars using DBSCAN. Jerabkova et al. (2021) used the compact convergent point method to recover 800 pc long tidal tails of the Hyades star cluster using Gaia DR2 and EDR3.
@@
Using Gaia DR3, Hunt & Reffert (2023) applied the HDBSCAN algorithm to recover 7167 clusters, also including tidal tails for some clusters. However, Hunt & Reffert (2024) derived completeness-corrected photometric masses for 6956 clusters from the earlier work and used these masses to compute their Jacobi radius ($r_{J}$). Kos (2024) simulated cluster dissolution to estimate the likelihood of star positions, comparing these with Gaia DR3 data to refine membership probabilities. Recently, Risbud et al. (2025) used the convergent point method to identify tidal tails in 19 open clusters within 400 pc.
@@
The previous methods often struggled with limitations such as poor astrometric precision, model dependencies, or applicability limited to nearby clusters. These constraints significantly reduce the number of clusters that can be studied in detail. To address this, we designed a method that uses optimized clustering and kinematic analysis to identify tidal tails even in distant systems. In this work, we have used the Gaia DR3 dataset to identify and investigate tidal tails in five open clusters: BH 164, NGC 2354, Alessi 2, NGC 2281, and M67. We tuned the algorithms to increase the detected extent of tidal tails, incorporating radial velocity (RV) filtering (where available), color-magnitude diagram (CMD) matching for age and composition consistency, and an unsupervised approach using PCA to determine the principal axis of elongation, followed by DBSCAN for tail detection. The identified tails were validated using the Jacobi limit to confirm their tidal origin. Projection effects were taken into account in the tangential plane, providing a clearer spatial distribution of the tails. We detected a greater number of stars in the four clusters and identified longer tidal tails in three of them compared to previous studies. Additionally, we obtained a more accurate estimate of the projected tail extent for M67, whose orientation aligns well with the expected orbit, consistent with N-body simulations. Furthermore, rotation was studied in the M67 and NGC 2281 clusters and tails. The residual velocity method, adopted from Hao et al. (2022, 2024), was previously used exclusively for member stars to confirm internal rotation in clusters such as Praesepe, Pleiades, Alpha Persei, and Hyades. In this study, we applied it to the cluster and its tails to analyze their rotational patterns.
@@
Section 2 outlines the data preparation, identification of the core cluster, and also explains the N-body simulations used for reference. Section 3 presents the methods used to detect tidal tails. In Sect. 4, we present a detailed analysis utilizing proper motion vector analysis, orbit integration, luminosity functions, and the binary fractions to explore the cluster and tail dynamics. Section 5 presents our tidal tail findings and discusses the results. In Sect. 6, the rotational aspect of M67 and NGC 2281 is explored using the residual motion of stars to study the presence of three-dimensional rotation. Finally, we discuss and summarize our research in Sects. 7 and 8, respectively.
@@ 2.2 Gaia DR3
The latest Gaia Data Release 3 (DR3; Gaia Collaboration et al. 2023) provides a substantial update over Gaia DR2, contains a total of 34 months of observational data, and features major advancements in data processing techniques. Gaia DR3 catalogs 1.47 billion sources with 5- or 6-parameter astrometric solutions, achieving a 30% improvement in parallax precision and nearly doubling the accuracy of proper motion measurements. These enhancements significantly improve the visibility of open clusters within the Gaia dataset, particularly in proper motion diagrams. The signal-to-noise ratio (S/N) for identifying distant open clusters has increased by approximately four (Hunt & Reffert, 2023), benefiting from a more compact Gaussian distribution of cluster stars for those with proper motion dispersions now within the reduced Gaia error margins.
@@
  • BP–RP (in mag): color index from Gaia blue and red photometers. Typical errors: $\lesssim$0.01 mag (bright), $\approx$0.05 mag (faint).
@@
  • RV (in km s-1): line-of-sight velocity, available for bright stars ($G\lesssim 15$). Typical errors: $\approx$0.2–5 km s-1.
Figure 1: Spatial distribution, proper motion, and parallax of the filtered sample (light coral) after filtering and the cluster members (blue) after applying DBSCAN for M67.
@@ 2.3 Data filtering and preparation
  • Proper motion filtering: pmra_min $\leq$ pmra $\leq$ pmra_max and  pmdec_min $\leq$ pmdec $\leq$ pmdec_max.
    For example, for M67, the ranges used were: -12.5 mas yr-1 $\leq$ pmra $\leq$ -9.5   mas yr-1,  -4 mas yr-1 $\leq$ pmdec $\leq$ -1 mas yr-1. These bounds were chosen based on the literature values (Gao, 2020a) and by visually inspecting the star distribution in proper motion space.
  • Error filtering: Stars with proper motion errors <0.5<0.5 mas yr-1 were retained. We limited the parallax using $\varpi>0$ and $\sigma_{\varpi}/\varpi<0.1$. Stars passing these filters were labeled as the filtered sample. We used relative parallax error to maintain consistent distance precision across clusters, although this imposes stricter absolute error thresholds on distant clusters, potentially biasing against the inclusion of faint members. Additionally, unresolved binaries may introduce astrometric errors not explicitly accounted for in this analysis. The spatial, proper motion, and parallax distributions for M67 are highlighted in light coral in Fig. 1.
  • Data preparation: For core cluster identification, we used the astrometric features: ra, dec, pmra, pmdec, and distance. We used distance instead of parallax to linearize spatial separations, enabling meaningful Euclidean distances in the clustering feature space. Since we filtered stars to have a relative parallax error <0.1<0.1, these distances are considered reliable. Stars with missing (NaN) values in any of these features were excluded. These core features were then normalized to zero mean and unit variance to ensure each contributed equally to the clustering process.
@@ 2.4 DBSCAN clustering for core cluster identification
Density-Based Spatial Clustering of Applications with Noise (Ester et al., 1996; Pedregosa et al., 2011) is a density-based clustering algorithm that identifies groups of closely packed points while treating outliers as noise. Unlike methods such as $k$-means, DBSCAN does not require specifying the number of clusters in advance and can discover clusters of arbitrary shape, making it well-suited for astrophysical data where the boundaries of a star cluster may not be sharply defined. It gives clear-cut clusters without soft membership boundaries.
For a data point to be considered as a part of a cluster, it must have a minimum number of neighboring data points within a specified distance, $\epsilon$. Key Parameters of DBSCAN:
@@
The DBSCAN algorithm identifies core points, border points, and noise points. Core points have sufficient neighbors within $\epsilon$; border points are neighbors of core points but are not core themselves; noise points are neither. In this work, DBSCAN was applied to the scaled 5D feature (ra, dec, pmra, pmdec, distance) space of the filtered sample to identify a dense overdensity corresponding to the core cluster population. For each cluster, the parameters ($\epsilon$, min_neighbors) were tuned to find the combination that resulted in the most spatially complete and centrally concentrated core in ra–dec space, while minimizing contamination. In M67, this method isolated 1037 cluster members (Nmemb), whose spatial, proper motion, and parallax distributions are also shown in blue in Fig. 1.
@@ 3 Tidal tail detection
Once core members were identified, we searched for tidal tail stars among noncluster filtered sample stars within 500 pc of the cluster’s mean distance. The approach combined RV filtering, CMD matching, and machine learning with scikit-learn (Pedregosa et al., 2011), including k-d tree (k-dimensional tree), PCA, and DBSCAN. Tidal stars are expected to share RVs and CMD features with the cluster, align along the cluster’s elongation axis, and lie beyond their Jacobi limit.
Table 1: Core cluster parameters of the selected open clusters. [^1] Col. 1: Cluster name; Cols. 2–3: Equatorial coordinates; Cols. 4–5: Galactic lon-lat; Col. 6: Age from Bossini et al. (2019) and Cantat-Gaudin et al. (2020); Cols. 7-10: Proper motions, mean parallax, and mean distance of identified members; Col. 11: Mean radial velocity of identified core cluster members; Col. 12: Number of identified core cluster members.
Name ra dec $l$ $b$ Age pmra pmdec $\varpi$ Dist RV $N_{\rm memb}$
@@ 3.1 Radial velocity filtering
After defining the core cluster, we applied an RV filter to further refine the potential tail candidates. Specifically, we selected stars within the 500 pc distance range whose RVs fall within a range of $\pm 10$ km s-1 of the core cluster’s mean RV:
$\texttt{RV}_{\text{tail}}=\texttt{RV}_{\text{core}}\pm 10\,\text{km s${}^{-1}$}.$ (1)
This $\pm 10$ km s-1 threshold serves as a loose filtering criterion to remove clear RV outliers that are inconsistent with the cluster’s dynamics. It is broad enough to account for small projection-induced velocity gradients [^2] Projection effects are corrected after the tidal tail detection for further analysis; see Sect. 4.1., which arise from geometric effects over large angular separations. This window was chosen to preserve plausible tail candidates while excluding clear outliers. Since Gaia DR3 provides RVs only for a subset of stars, this filtering step is applied only where RV data is available. For stars lacking RV data, we rely on CMD-based filtering to evaluate their consistency with the cluster’s stellar population.
3.2 Color-magnitude diagram matching
In addition to RV, we used the color (BP-RP) versus apparent G magnitude CMD to refine our identification of tidal tail stars. In the CMD space, tidal tail stars are expected to closely resemble the core cluster members, sharing similar age, metallicity, and evolutionary history. The CMD provides a visual representation of these properties and allows us to distinguish between stars that formed together and those that are unrelated to background or foreground stars.
To efficiently process large datasets such as Gaia DR3, we employed a k-d tree structure for performing the CMD cut. This algorithm enables fast nearest-neighbor searches in high-dimensional space and was used to identify stars whose CMD values closely match those of the core cluster members. The tree partitions the data recursively by splitting on dimensions such as color and magnitude. A distance threshold of 0.05 in the CMD space was applied to select potential tidal tail members, ensuring photometric similarity with the cluster core. The CMD distance is defined as
$d_{\text{CMD}}=\sqrt{(G-G_{\text{core}})^{2}+\left[(BP-RP)-(BP_{\text{core}}-RP_{\text{core}})\right]^{2}}.$ (2)